Find the sum of the first eight terms of each of these geometric series.
a) 729 + 243 + 81 + 27 + …
b) 2 - 6 + 18 - 54 + …

Answer by Boreal(15235)

s= a (1-r^(n+1))/(1-r)
the first term has a=729 and r=(1/3), and n+1=8, since first term is ar^0
so s=(729) (1-(1/3)^8)/(1-(1/3))
=729 (1.49977) don't round
=1093 (1/3)
729+243+81+27+9+3+1+(1/3)=1093 (1/3)
--
this is r=(-3)
s=2(1-(-3)^8/(1-(-3))
=2*(1-6561)/4
=-3280
2-6+18-54+162-486+1458-4374=-3280